# Can Authorship Attribution Models Distinguish Speakers in Speech Transcripts?

### v2

@@ 4.2 Experiment 2: Transcription style
To test the impact of transcription style, we ran Exp 1 (with the same verification trials) on both the BBN encoding, with punctuation and capitalization, and the LDC encoding, with limited or none. Overall, we find that transcription style can have a surprisingly large impact on performance.
@@
Comparing the leftmost column (‘base’) of the left (BBN) and middle (LDC) sections of Table 2, we see that TF-IDF, PANgrams, and SBERT are similar for both encodings. A lack of difference for these models is expected since semantic content is similar across embeddings. AdHominem shows some improvement from BBN to LDC, but there are significant increases for CISR and especially LUAR with the LDC encoding. The ability of LUAR to perform well out-of-the-box, and to perform significantly better on the normalized transcription style in general, suggests that LUAR does not rely on ‘superficial’ prescriptive textual features. The fact that model performance does not degrade on normalized data, even improving in some cases, is a promising sign for potential applications to automatically transcribed speech, which often lacks such features.
@@
Since LUAR showed the biggest difference between encodings, we ran a LUAR model pre-trained on the same subset of the Reddit dataset as before, this time normalized to look like the LDC encoding. The normalization included lowercasing and removing all HTML special entities, hyperlinks, emoticons, and punctuation except apostrophes and hyphens between letters. This model is called LUARnorm and its out-of-the-box performance is in Table 3. LUAR and LUARnorm achieve similar performance.
Since the speech transcripts contain bracketed non-speech sounds and annotators’ hypothesized transcriptions, the normalized Reddit dataset and LDC encoding are still not exactly equivalent. Thus, we also further normalized the LDC encoding data, creating a NormLDC encoding by removing the brackets and double parentheses. This new encoding now more closely resembles the normalized Reddit, leaving only differences between text and speech characteristics. Both LUAR and LUARnorm perform best on the NormLDC encoding out of all three encodings and continue to perform roughly similarly. These results suggest that a lack of capitalization and punctuation (LDC), along with the removal of transcript-specific annotations (NormLDC), seem most influential for improving performance, but pre-training on normalized text data (LUARnorm) does not significantly impact performance.
@@ 4.3 Experiment 3: Topic
To test the effect of topic, after running Exp 1 on the ‘base’ dataset, we ran it on the ‘hard’ and ‘harder’ datasets. Since performance of both LUAR (and LUARnorm) was highest on the NormLDC encoding previously, we included it here as well. The full results across all settings are shown in Table 2.
@@
Across all three encodings, the ‘hard’ dataset had lower AUC scores than the ‘base’ dataset, and the ‘harder’ dataset lower than the ‘hard’ dataset. In particular, TF-IDF and AdHominem decrease consistently to around chance. Even PANgrams, one of the best performers on the ‘base’ level, degrades significantly, suggesting that simple -gram approaches are incapable of capturing speaker stylistic features and, as a result, fail under topic shift. SBERT was the most severely affected by the topic manipulation, with performance on the hardest dataset well below chance. As we expect SBERT to exploit content differences that we successively remove with this manipulation, this result is not surprising. CISR also had a decrease in performance but to a smaller extent, likely because its training involves specific attempts at controlling for topic by using subreddits as proxies for topic in the Reddit training data (Wegmann et al., 2022).
For LUAR, we see ~10% performance loss between the ‘base’ and ‘hard’ conditions, and ~15% between the ‘hard’ and ‘harder’ conditions, to near chance. We propose two potential explanations for the greater difference between the ‘hard’ and ‘harder’ conditions. First, the ‘hard’ condition does not fully accomplish its topic manipulation as proxies for topic diverge from linguistic definitions (e.g. ‘summer plans’ may include discussions of both vacations and temp jobs, with little semantic overlap between these topics). With the ‘harder’ dataset, in which each speaker in the trial discusses largely the same topic(s) and subtopic(s) in the same order given that they represent each side of the same conversation, we expect a higher degree of topic identity.
@@
Overall, since topic shifts are expected to some degree in many applications of speaker verification, these results suggest that methods developed for written text are inadequate out-of-the-box for verification of transcribed speech. CISR is a promising approach for being more resilient to topic control, but performs worse compared to other models on the ‘base’ and ‘hard’ levels.
@@ 4.4 Experiment 4: Fine-tuning
TF-IDF | 0.500 | 0.519 | 0.491 | 0.493 | 0.524 | 0.521 | 0.532 | 0.544 | 0.523 |
PANgrams | 0.763 | 0.628 | 0.419 | 0.764 | 0.623 | 0.412 | 0.765 | 0.628 | 0.417 |
AdHominem | 0.586 | 0.578 | 0.541 | 0.594 | 0.576 | 0.546 | 0.542 | 0.558 | 0.586 |
SBERT | 0.694 | 0.650 | 0.632 | 0.589 | 0.830 | 0.835 | 0.530 | 0.818 | 0.935 |
CISR | 0.722 | 0.696 | 0.744 | 0.660 | 0.642 | 0.690 | 0.674 | 0.651 | 0.781 |
LUAR | 0.844 | 0.818 | 0.753 | 0.798 | 0.872 | 0.818 | 0.694 | 0.820 | 0.895 |
To see whether fine-tuning on speech transcripts can improve transfer from authorship models to speech transcripts, we train a multilayer perceptron (MLP) classifier on the training set verification trials across the models.101010Model selection was performed based on validation performance. We found that other fine-tuning approaches, such as linear models, performed worse on validation data. PANgrams did not use a MLP but instead directly used the most recent PAN baseline code (Stamatatos et al., 2023) Models are fine-tuned on the training set for each difficulty level and then evaluated on all test sets, e.g. train on ‘base’ and evaluate this model on ‘base’, ‘hard’, and ‘harder’. Table 4 gives these results for the LDC encoding, but the results for the other encodings followed the same pattern.
@@
The SBERT model achieves the best performance in the ‘harder’-‘harder’ condition (0.935). Unlike SBERT, which was not able to overcome the reduction in content differences in the ‘harder’ setting, SBERT seems to have ‘learned’ to take advantage of the content similarity. Returning to rates of noun overlap in Table 1, there is almost double the noun overlap between transcripts of different speakers in the same conversation (negative trials in the ‘harder’ setting) than between transcripts of the same speaker in conversations on different topics (positive trials in the ‘harder’ setting). One hypothesis for SBERT’s higher performance, then, is that it used the rate of noun overlap as a negative indicator of whether the speaker is the same or not, with higher noun overlap rates indicating different speakers and lower rates indicating the same speaker. In other words, SBERT used a shortcut that may have worked well in this experimental setup, but likely would not work well in other setups, which is corroborated by its poorer performance on the ‘harder’-‘base’ and ‘harder’-‘hard’ settings.
LUAR has the next best performance after SBERT, also in the ‘harder’-‘harder’ condition, and shows a strong correspondence between distribution matching. CISR, though, does not show this same pattern. Again, as its training attempts to control for topic, its performance is less affected by our topic manipulation, and fine-tuning on particular settings does not show as significant of a difference. AdHominem could be similar, though its performance in general is much lower.
Unlike the out-of-the-box models, the fine-tuned models are able, to an extent, to overcome challenges, such as less-than-perfect topic control and speaker style accommodation due to their exposure to trials of the same, or similar, difficulty level in training, to encode identifying stylistic features of speakers beyond the conversation topic. However, there is no general-purpose model that works well across difficulty levels; models work best when they are trained and evaluated on data from the same distribution. These models have potential for further improvement with more specialized tuning, which is a direction for future work.
@@ 4.5 Experiment 5: Pre-training domain
LUAR | 0.837 | 0.722 | 0.543 |
LUARnorm | 0.831 | 0.704 | 0.524 |
PreTrain-BBN | 0.887 | 0.709 | 0.505 |
PreTrain-LDC | 0.906 | 0.699 | 0.418 |
LUAR | 0.844 | 0.869 | 0.876 |
LUARnorm | 0.864 | 0.875 | 0.907 |
PreTrain-BBN | 0.906 | 0.909 | 0.952 |
PreTrain-LDC | 0.909 | 0.921 | 0.935 |
@@
PreTrain-BBN and PreTrain-LDC are two separate instantiations of LUAR that were pre-trained on the full training set of speech transcripts, without refining by difficulty level, for the BBN and LDC encoding, respectively. Out-of-the-box, these models followed the previously described pipeline of being evaluated using cosine similarity on the validation (and later test) set. In the fine-tuned case, again following the same protocol as before, an MLP classifier was trained on the training set verification trials (i.e. the same training data as seen in pre-training, but this time as trials of a particular difficulty) and evaluated using AUC score on the validation (and later test) set. Since training and evaluating on the same difficulty performed best in Exp 4, we restrict the experiment to the train-test distribution match condition across difficulty levels. Performance was best on the NormLDC encoding, so Table 5 focuses on these results.
As expected, compared to the other LUAR models, PreTrain-BBN and PreTrain-LDC perform best across all three difficulties and achieve the highest performance of any model on the ‘harder’ level. On the ‘base’ level, PreTrain-BBN and PreTrain-LDC are fairly close seconds to their fine-tuned counterparts; however, in the ‘hard’ and ‘harder’ levels, PreTrain-BBN’s and PreTrain-LDC’s performance decreases significantly. This drop-off is most likely due to the train-test mismatch between pre-training on all training transcripts and evaluating only on verification trials of a particular difficulty level. The fine-tuned models suggest, though, through their increased performance across levels, that during pre-training, LUAR encodes speech-specific features that the MLP can avail of.
@@ 4.6 Experiment 6: Varying input size
Our final experiment tested the impact observing more data has on attribution by varying the number of utterances used in each verification pair. We ran Exp 1 and Exp 4 (with training and evaluation distributions matched) on trials of transcripts containing incrementally more utterances, ranging from the first 25 per speaker to the full transcript. Speakers averaged ~95 utterances per transcript (after the first 5 were removed) and the longest had ~200 utterances. We chose the four best performing models for this experiment but only one LUAR (the standard instantiation for better comparability): PANgrams, SBERT, CISR, and LUAR. Since performance across models was not consistently better on the NormLDC encoding, we ran this experiment only on the BBN and LDC encodings, and, given similar resulting patterns across these two, we present graphs only for the LDC encoding in Figure 2. These display the AUC score performance for each out-of-the-box (left column) and fine-tuned model (right column) on transcripts of length 25, 75, 135, and the full number of utterances per speaker in each pair. Each row represents an increase in difficulty level.
@@
Across both out-of-the-box and fine-tuned models, all increase in performance from 25 to 75 utterances, beginning to plateau after 75, and some reach slightly higher performance by 135 utterances. Only 22% of the LDC test transcripts are less than 75 utterances long and 67% are less than 100 utterances, so most transcripts have not concluded at 75, yet the end of the transcript does not seem to add much extra information given performance plateaus. Future work should divide each transcript non-cumulatively, comparing e.g. beginning, middle, and end, to help determine which sections most reveal speaker style.
@@ 5 Discussion
The primary goal of this work was to provide a proof of concept and establish baseline performance of text-based authorship models on speech transcripts. In so doing, we discovered the following answers to our research questions. First, despite the modality difference, we found that off-the-shelf textual authorship models, such as PANgrams and LUAR, transfer surprisingly well to speech transcripts, unless we control for topic, in which case all models’ performance drops drastically. This finding contradicts previous work that did not rigorously control for topic, suggesting that model performance may have resulted from spurious correlations with topic rather than an ability to distinguish speakers.
Transcription style also impacts performance, with most models performing best on LDC transcripts normalized to remove capitalization and punctuation. Further normalization to remove speech transcript-specific annotations, such as bracketed non-speech sounds and annotators’ hypothesized transcriptions (NormLDC), hurt performance for many models, however. Adding superficial prescriptive textual features to transcripts is thus perhaps an unneeded processing step, but maintaining a distinction between annotations and regular speech is.
We found that fine-tuning on speech transcripts significantly improves performance for most neural models, with the best performance achieved when the training and evaluation data are drawn from the same difficulty-level distributions, specifically in the ‘harder’ condition where negative pairs are drawn from the same conversation. The ‘base’ setting, though, represents a complete sample of pairwise verification trials without any artificial subsampling; therefore, while it is our easiest setting, it also represents a possibly more realistic setting than ‘hard’ and ‘harder’ cases. Choice of model should thus be determined by the particularities of the available data and the specific application. Separately, we find that additionally pre-training the model on speech transcripts can further improve performance. Finally, performance across all models plateaus after 75 utterances, despite most transcripts containing at least 20 more utterances, suggesting that conversation at the end of the call does not contribute much extra identifying information of the speaker.
@@
We note that in our ‘harder’ dataset, accommodation may play a role in the results in addition to topic control, but do not tease apart the relative impact of this and other factors. Future work should attempt to quantify the amount of accommodation that occurs between speakers in the same call. Comparing performance on different partitions of the transcript, such as truncating each transcript to the beginning, middle, or end third, could help further inform which stages of conversation are most revealing of speaker style. An eventual extension might also look at the extent to which topics change over the course of conversations with specified discussion subjects, though a qualitative evaluation of some of the corpus indicated less evolution than expected.

### v3

@@ 4.2 Experiment 2: Transcription style
To test the impact of transcription style, we ran Experiment 1 (with the same verification trials) on both the BBN encoding, with punctuation and capitalization, and the LDC encoding, with limited or none. Overall, we find that transcription style can have a surprisingly large impact on performance.
@@
Comparing the leftmost column (‘base’) of the left (BBN) and middle (LDC) sections of Table 2, we see that TF-IDFo, PANgramso, and SBERTo are similar for both encodings. A lack of difference for these models is expected since semantic content is similar across embeddings. AdHominemo shows some improvement from BBN to LDC, but there are significant increases for CISRo and especially LUARo with the LDC encoding. The ability of LUARo to perform well out-of-the-box, and to perform significantly better on the normalized transcription style in general, suggests that LUAR does not rely on ‘superficial’ prescriptive textual features. The fact that model performance does not degrade on normalized data, even improving in some cases, is a promising sign for potential applications to automatically transcribed speech, which often lacks such features.
@@
Since LUAR showed the biggest difference between encodings, we ran a LUAR model pre-trained on the same subset of the Reddit dataset as before, this time normalized to look like the LDC encoding. The normalization included lowercasing and removing all HTML special entities, hyperlinks, emoticons, and punctuation except apostrophes and hyphens between letters. This model is called LUARnorm and its out-of-the-box performance on the ‘base’ level is in Table 3. LUARo and LUARnormo achieve similar performance. (The performance of both models across all encodings and difficulties is shown in Table 9 (AUC) and Table 10 (EER).)
Since the speech transcripts contain bracketed non-speech sounds and annotators’ hypothesized transcriptions, the normalized Reddit dataset and LDC encoding are still not exactly equivalent. Thus, we also further normalized the LDC encoding data, creating a NormLDC encoding by removing the brackets and double parentheses. This new encoding now more closely resembles the normalized Reddit, leaving only differences between text and speech characteristics. Both LUARo and LUARnormo perform best on the NormLDC encoding out of all three encodings and continue to perform roughly similarly. These results suggest that a lack of capitalization and punctuation (LDC), along with the removal of transcript-specific annotations (NormLDC), seem most influential for improving performance, but pre-training on normalized text data (LUARnorm) does not significantly impact performance.
@@ 4.3 Experiment 3: Topic
To test the effect of topic, after running Experiment 1 on the ‘base’ dataset, we ran it on the ‘hard’ and ‘harder’ datasets. Since performance of both LUARo (and LUARnormo) was highest on the NormLDC encoding previously, we included it here as well. The full AUC results across all settings are shown in Table 2 and the corresponding EER results are in Table 8.
@@
Across all three encodings, the ‘hard’ dataset had lower AUC scores than the ‘base’ dataset, and the ‘harder’ dataset lower than the ‘hard’ dataset. In particular, TF-IDFo and AdHominemo decrease consistently to around chance. Even PANgramso, one of the best performers on the ‘base’ level, degrades significantly, suggesting that simple -gram approaches are incapable of capturing speaker stylistic features and, as a result, fail under topic shift. SBERTo was the most severely affected by the topic manipulation, with performance on the hardest dataset well below chance. As we expect SBERTo to exploit content differences that we successively remove with this manipulation, this result is not surprising. CISRo also had a decrease in performance but to a smaller extent, likely because its training involves specific attempts at controlling for topic by using subreddits as proxies for topic in the Reddit training data (Wegmann et al., 2022).
For LUARo, we see ~10% performance loss between the ‘base’ and ‘hard’ conditions, and ~15% between the ‘hard’ and ‘harder’ conditions, to near chance. We propose two potential explanations for the greater difference between the ‘hard’ and ‘harder’ conditions. First, the ‘hard’ condition does not fully accomplish its topic manipulation as proxies for topic diverge from linguistic definitions of thematic topic, discourse topic, and the like. As two examples, ‘summer plans’ may include discussions of both vacations and temp jobs, with little semantic overlap between these topics, and a range of subtopics may be discussed throughout each conversation. With the ‘harder’ dataset, in which each speaker in the trial discusses largely the same topic(s) and subtopic(s) in the same order given that they represent each side of the same conversation, we expect a higher degree of topic identity.
@@
Overall, since topic shifts are expected to some degree in many applications of speaker verification, these results suggest that methods developed for written text are inadequate out-of-the-box for verification of transcribed speech. CISRo is a promising approach for being more resilient to topic control, but performs worse compared to other models on the ‘base’ and ‘hard’ levels.
@@ 4.4 Experiment 4: Fine-tuning
| TF-IDFft | 0.500 | 0.519 | 0.491 | 0.493 | 0.524 | 0.521 | 0.532 | 0.544 | 0.523 |
| PANgramsft | 0.763 | 0.628 | 0.419 | 0.764 | 0.623 | 0.412 | 0.765 | 0.628 | 0.417 |
| AdHominemft | 0.586 | 0.578 | 0.541 | 0.594 | 0.576 | 0.546 | 0.542 | 0.558 | 0.586 |
| SBERTft | 0.694 | 0.650 | 0.632 | 0.589 | 0.830 | 0.835 | 0.530 | 0.818 | 0.935 |
| CISRft | 0.722 | 0.696 | 0.744 | 0.660 | 0.642 | 0.690 | 0.674 | 0.651 | 0.781 |
| LUARft | 0.844 | 0.818 | 0.753 | 0.798 | 0.872 | 0.818 | 0.694 | 0.820 | 0.894 |
To see whether fine-tuning on speech transcripts can improve transfer from authorship models to speech transcripts, we fit a multilayer perceptron (MLP) classifier to the concatenated embeddings, from each model, of each trial in the training set verification trials.131313Model selection was performed based on validation performance. We found that other fine-tuning approaches, such as linear models, performed worse on validation data. Hyperparameter optimization experiments found that using the Adam solver with 800 maximum iterations worked best overall across models on the validation trials. We kept all other default values and used a random state of 1. Since the available number of trials is limited based on the size of the corpus, we are wary of overfitting the data and thus fit a transformation of the (fixed) embedding from each model rather than fine-tune the whole model. We then evaluate the classifier on its probability predictions of the concatenated embeddings of the test trials. To account for variation, we bootstrap the AUC score over 1000 random resamples of the test trials. We use this procedure for all models except PANgramsft, which directly uses the PAN-provided code to train and calculate cosine similarity within trials (Stamatatos et al., 2023), but we additionally bootstrap its AUC score over 1000 random test trial resamples to produce a more robust evaluation.
The classifier is fine-tuned on the training set for each difficulty level and then evaluated on all test sets, e.g. train on ‘base’ and evaluate this model on ‘base’, ‘hard’, and ‘harder’. Table 4 gives these results for the LDC encoding, which produced higher scores overall than the BBN encoding. Table 7 shows the results for the BBN encoding; EER results for the train-test match setting across the BBN, LDC, and NormLDC encodings are provided in Table 8.
@@
The SBERTft model achieves the best performance in the ‘harder’-‘harder’ condition (0.935). Unlike SBERTo, which was not able to overcome the reduction in content differences in the ‘harder’ setting, SBERTft seems to have ‘learned’ to take advantage of the content similarity. Returning to rates of noun overlap in Table 1, there is almost double the noun overlap between transcripts of different speakers in the same conversation (negative trials in the ‘harder’ setting) than between transcripts of the same speaker in conversations on different topics (positive trials in the ‘harder’ setting). One hypothesis for SBERTft’s higher performance, then, is that it used the rate of noun overlap as a negative indicator of whether the speaker is the same or not, with higher noun overlap rates indicating different speakers and lower rates indicating the same speaker. In other words, SBERTft used a shortcut that may have worked well in this experimental setup, but likely would not work well in other setups, which is corroborated by its poorer performance on the ‘harder’-‘base’ and ‘harder’-‘hard’ settings.
LUARft has the next best performance after SBERTft, also in the ‘harder’-‘harder’ condition, and shows a strong correspondence between distribution matching. CISRft, though, does not show this same pattern. Again, as its training attempts to control for topic, its performance is less affected by our topic manipulation, and fine-tuning on particular settings does not show as significant of a difference. AdHominemft could be similar, though its performance in general is much lower.
Unlike the out-of-the-box models, the fine-tuned models can, to an extent, overcome challenges, such as less-than-perfect topic control and speaker style accommodation; exposure to trials of the same, or similar, difficulty level in training enables them to encode identifying stylistic features of speakers beyond the conversation topic. However, there is no general-purpose model that works well across difficulty levels; models work best when they are trained and evaluated on data from the same distribution. These models have potential for further improvement with more specialized tuning, which is a direction for future work.
@@ 4.5 Experiment 5: Pre-training domain
| LUARo | 0.837 | 0.722 | 0.543 |
| LUARnormo | 0.831 | 0.704 | 0.524 |
| PreTrain-BBNo | 0.887 | 0.709 | 0.505 |
| PreTrain-LDCo | 0.906 | 0.699 | 0.418 |
| LUARft | 0.844 | 0.869 | 0.876 |
| LUARnormft | 0.864 | 0.875 | 0.907 |
| PreTrain-BBNft | 0.906 | 0.909 | 0.952 |
| PreTrain-LDCft | 0.909 | 0.921 | 0.935 |
@@
PreTrain-BBN and PreTrain-LDC are two separate instantiations of LUAR that were pre-trained on the full training set of speech transcripts, without refining by difficulty level, for the BBN and LDC encoding, respectively. Out-of-the-box, these models followed the previously described pipeline of being evaluated using cosine similarity on the validation (and later test) set. In the fine-tuned case, again following the same protocol as before, an MLP classifier was trained on the training set verification trials (i.e. the same training data as seen in pre-training, but this time as trials of a particular difficulty) and evaluated using bootstrapped AUC score on the validation (and later test) set. Since training and evaluating on the same difficulty performed best in Experiment 4, we restrict the experiment to the train-test distribution match condition across difficulty levels. Performance was best on the NormLDC encoding, so Table 5 focuses on these results, but the results for all encodings are shown in Table 9 (AUC) and Table 10 (EER).
As expected, compared to the other LUAR models, PreTrain-BBNft and PreTrain-LDCft perform best across all three difficulties and achieve the highest performance of any model on the ‘harder’ level. On the ‘base’ level, PreTrain-BBNo and PreTrain-LDCo are fairly close seconds to their fine-tuned counterparts; however, in the ‘hard’ and ‘harder’ levels, PreTrain-BBNo’s and PreTrain-LDCo’s performance decreases significantly. This drop-off is most likely due to the train-test mismatch between pre-training on all training transcripts and evaluating only on verification trials of a particular difficulty level. The fine-tuned models suggest, though, through their increased performance across levels, that during pre-training, LUAR encodes speech-specific features that the MLP can avail of.
@@ 4.6 Experiment 6: Varying input size
Our final experiment tested the impact that observing more data has on attribution by varying the number of utterances used in each verification pair. We ran Experiment 1 and Experiment 4 (with training and evaluation distributions matched) on trials of transcripts containing incrementally more utterances, ranging from the first 25 per speaker to the full transcript. Speakers averaged ~95 utterances per transcript (after the first 5 were removed) and the longest had ~200 utterances. We chose the four best performing models for this experiment but only one LUAR (the standard instantiation for better comparability): PANgrams, SBERT, CISR, and LUAR. Since performance across models was not consistently better on the NormLDC encoding, we ran this experiment only on the BBN and LDC encodings. The graphs for the LDC encoding are in Figure 2 and those for the BBN encoding are in Figure 3. These display the AUC score performance for each out-of-the-box (left column) and fine-tuned model (right column) on transcripts of length 25, 75, 135, and the full number of utterances per speaker in each pair. Each row represents an increase in difficulty level.
@@
Across both out-of-the-box and fine-tuned models, all increase in performance from 25 to 75 utterances, begin to plateau after 75, and some reach slightly higher performance by 135 utterances. Only 22% of the LDC test transcripts are less than 75 utterances long and 67% are less than 100 utterances, so most transcripts have not concluded at 75.
| LDC Encoding | ||||||
| Base | Hard | Harder | ||||
| Model | first 50 | last 50 | first 50 | last 50 | first 50 | last 50 |
| PANgramso | 0.741 | 0.757 | 0.620 | 0.683 | 0.461 | 0.519 |
| SBERTo | 0.648 | 0.675 | 0.467 | 0.654 | 0.244 | 0.309 |
| CISRo | 0.747 | 0.646 | 0.682 | 0.664 | 0.637 | 0.600 |
| LUARo | 0.723 | 0.782 | 0.682 | 0.767 | 0.495 | 0.441 |
To help determine if different parts of the transcript contribute more speaker information than others, we also ran a preliminary experiment attributing transcripts using only the beginning or the end of the transcript. Specifically, we restrict to transcripts having at least 100 utterances and create two further evaluation datasets. The first is similar to our existing experiments but restricted to the first 50 utterances to produce the speaker representation. The second instead takes the last 50 utterances to produce the speaker representation. Then we construct three experiments for our ‘base’, ‘hard’, and ‘harder’ settings using the same protocol as elsewhere in the paper. Since there were significantly fewer trials (~100 per difficulty) after requiring each speaker per trial to have at least 100 utterances, these results should be compared relatively to each other, but not to other results in the paper. Table 6 reports these results for the same four best performing out-of-the-box models on the LDC encoding.
We make the following observations. First, we see that the baseline model PANgramso and the semantic model SBERTo consistently benefit from using the last 50 utterances across all difficulty levels. LUARo, which performs better in the ‘base’ and ‘hard’ settings, also benefits from using the last 50 utterances over the first 50, except in the ‘harder’ setting. The CISRo style representation, however, exhibits the opposite trend as PANgramso and SBERTo, being consistently worse when using the last 50 utterances. Overall, further study is needed to understand these differences in performance, although we may hypothesize that accommodation, which is likely to be evident later in a conversation, is playing a role in explaining these results, particularly for CISRo.
@@ 5 Discussion
The primary goal of this work was to provide a proof of concept and establish baseline performance of text-based authorship models on speech transcripts. We focused on gold standard, human-transcribed transcripts as a starting point to find an upper bound on performance since this task has not previously been benchmarked. In so doing, we discovered the following answers to our research questions. First, despite the modality difference, we found that off-the-shelf textual authorship models, such as PANgrams and LUAR, transfer surprisingly well to speech transcripts, unless we control for topic, in which case all models’ performance drops drastically. This finding contradicts previous work that did not rigorously control for topic, suggesting that model performance may have resulted from spurious correlations with topic rather than an ability to distinguish speakers.
Transcription style also impacts performance, with most models performing best on the LDC transcripts normalized to remove capitalization and punctuation. Further normalization to remove speech transcript-specific annotations, such as brackets around non-speech sounds and double parentheses around annotators’ hypothesized transcriptions (NormLDC), hurt performance for many models, however. Adding superficial prescriptive textual features to transcripts is thus perhaps an unneeded processing step, but maintaining a distinction between annotations and regular speech is.
We found that fine-tuning on speech transcripts significantly improves performance for most neural models, with the best performance achieved when the training and evaluation data are drawn from the same difficulty-level distributions, specifically in the ‘harder’ condition where negative pairs are drawn from the same conversation. The ‘base’ setting, though, represents a complete sample of pairwise verification trials without any artificial subsampling; therefore, while it is our easiest setting, it also represents a possibly more realistic setting than ‘hard’ and ‘harder’ cases. Choice of model should thus be determined by the particularities of the available data and the specific application. Separately, we find that additionally pre-training the model on speech transcripts can further improve performance. Finally, performance across all models plateaus after 75 utterances, despite most transcripts containing at least 20 more utterances, but which section (beginning or end) of each call is most useful for speaker attribution differs by model.
@@
Our best results use fine-tuned author representations pre-trained on the same speech transcripts; future work should explore variations that might produce even better performance, such as pre-training a dedicated model on a larger and more diverse dataset of speech transcripts. To better understand how the models are performing, future work should conduct a qualitative analysis of the results, linguistically examining which trials the models predict correctly and incorrectly to find any consistencies across models as well as any features the models might be using to make their determinations.
We note that in our ‘harder’ dataset, accommodation may play a role in the results in addition to topic control, but do not tease apart the relative impact of this and other factors. Future work should attempt to quantify the amount of accommodation that occurs between speakers in the same call, which could also further inform which stages of conversation are most revealing of speaker style. An eventual extension might also look at the extent to which topics change over the course of conversations with specified discussion subjects, though a qualitative evaluation of some of the corpus indicated less evolution than expected.
@@
Ethical considerations
Our findings should be carefully interpreted before considering speaker attribution for any real-world applications. For instance, although fine-tuning a model in our ‘harder’ setting can significantly improve its performance on the same ‘harder’ condition, this is an artificial setting that is not generally representative of real-world distributions of speaker transcripts. In addition, we find that performance across all models significantly decreases with a more rigorous control for topic, indicating that it would be premature to apply these models if topic shifts or topic differences across samples are possible in the application domain. We acknowledge that further enhancements of the methods presented, such as a better accounting of topic, may be used to defeat speaker anonymization systems, but our results suggest that current methods are not yet robust enough to topic manipulations to have this capability. Regardless of topic, though, these models should not be applied in domains in which it is important to understand how and why an attribution decision is made; such models would not pass the Daubert standard (Daubert v. Merrell Dow Pharmaceuticals, Inc., 1993) for scientific evidence in the U.S. judicial context, for example.
Many of the models we used were trained on anonymized social media data that have implicit biases, such as imbalances in the prevalence of authors from certain demographic groups. As a result, attribution performance may vary based on the same latent demographic factors, which is an issue that needs further study. One remediation may be to control for demographics when training the neural representations and then ensure that within-group performance is consistent for all relevant factors. Finally, it is worth noting that population-level statistics do not currently exist to determine the extent to which people speak (or write) like one another—while it is tempting to think that we have unique patterns of speech based on the success of some attribution models, we still have no real understanding of how rare certain styles are.